# Globalization of Partial Actions of Ordered Groupoids on Rings

### v1

@@ 5.2. Partial Actions of Inverse Semigroups
Let $\alpha$ be a partial inverse semigroup action of $S$ on a ring $A$ such that $A_{e}$ is unital for every $e\in E(S)$. Then $\alpha$ has a minimal globalization $\beta$ if and only if $\alpha$ is unital. Moreover, $\beta$ is unique up to isomorphism.
Consider the definition of partial skew inverse semigroup ring given in [7], that is, $A\ltimes_{\alpha}S=\mathcal{L}/\mathcal{N}$, where $\alpha$ is a unital partial action of $S$ on a ring $A$, $\mathcal{L}=\bigoplus_{s\in S}A_{s}\delta_{s}$ with component-wise addition and product given by $(a_{s}\delta_{s})(b_{t}\delta_{t})=\alpha_{s}(\alpha_{s^{-1}}(a_{s})b_{t})\delta_{st}$ and $\mathcal{N}=\langle a_{s}\delta_{s}-a_{s}\delta_{t}:s\leq t\text{ and }a_{s}\in A_{s}\rangle$ is an ideal of $\mathcal{L}$. Then the same argument used in Theorem 5.2 also proves the following result.
@@
Let $\alpha$ be a unital partial inverse semigroup action of $S$ on a ring $A$. Assume that $E(S)$ is finite. Let $\beta$ be the minimal globalization of $\alpha$, global action of $S$ on a ring $B$. Then $A\ltimes_{\alpha}S$ and $B\ltimes_{\beta}S$ are Morita equivalent.

### v2

@@ 5.2. Partial Actions of Inverse Semigroups
Let $\alpha$ be a preunital partial action of an inverse semigroup $S$ on a ring $A$. Then $\alpha$ has a globalization $\beta$ if and only if $\alpha$ is unital. Moreover, $\beta$ is unique up to equivalences.
Proof.
Let $\gamma_{\alpha}$ be the inverse semigroup premorphism from $S$ to $\mathcal{I}(A)$ defined in Lemma 5.9. Let $\mathbb{G}(\gamma_{\alpha}):\mathbb{G}(S)\to\mathbb{G}(\mathcal{I}(A))$ be the inductive groupoid premorphism from the inductive groupoid $\mathbb{G}(S)$ to the inductive groupoid $\mathbb{G}(\mathcal{I}(A))$ associated to $\gamma_{\alpha}$ via the ESN Theorem. Since $\mathbb{G}(S)$ is inductive, it is pseudoassociative.
Notice that the functor $\mathbb{G}$ of the ESN Theorem is such that $F(S)=F(\mathbb{G}(S))$ and $F(\gamma_{\alpha})=F(\mathbb{G}(\gamma_{\alpha}))$ (considering the forgetful functor $F$ applied in the categories of inverse semigroups and of inductive groupoids, respectively). Hence, $\text{dom}(\mathbb{G}(\gamma_{\alpha})(s))=F(A_{s^{-1}})$ and $\text{Im}(\mathbb{G}(\gamma_{\alpha})(s))=F(A_{s})$. Therefore, defining $\omega=(A_{s},\omega_{s})_{s\in\mathbb{G}(S)}$ by $\omega_{s}:A_{s^{-1}}\to A_{s}$, with $\omega_{s}(a)=\mathbb{G}(\gamma_{\alpha})(s)(a)=\alpha_{s}(a)$, for all $a\in A_{s^{-1}}$, $s\in\mathbb{G}(S)$, we see that $\omega$ is a preunital strong P.O. action of $\mathbb{G}(S)$ on the ring $A$. Indeed, the conditions (P2) and (P3) are equivalent to the condition (i) of inductive groupoid premorphism. The condition (PO) corresponds to the condition (iii), while the condition (ii) of inductive groupoid premorphism is equivalent to [5, Lemma 1.1](i). Finally, [21, Definition 2.6] guarantees that (PS) holds, since all inductive groupoid premorphism $\psi$ satisfies $r(\psi(e|g))=\psi(e)\wedge r(\psi(g))$.
Now, we use Theorem 4.7 to conclude that $\omega$ admits minimal globalization if and only if is unital. Since $\alpha$ is unital if and only if $\omega$ is unital, $\alpha$ admits globalization if and only if $\omega$ has a minimal globalization. In this case, we obtain a global action $\eta=(B_{s},\eta_{s})_{s\in\mathbb{G}(S)}$ of $\mathbb{G}(S)$ on a ring $B$ that is a minimal globalization of $\omega$.
Next, analogously to the construction of $\gamma_{\alpha}$, define the inductive groupoid premorphism $\rho_{\eta}:\mathbb{G}(S)\to\mathbb{G}(\mathcal{I}(A))$ such that the function $\rho_{\eta}(s):F(B_{s^{-1}})\to F(B_{s})$ is given by $\rho_{\eta}(s)(b)=\eta_{s}(b)$, for all $b\in F(B_{s^{-1}})$. Notice that $\rho_{\eta}$ is indeed an inductive functor, because $\eta$ being global implies $B_{r(s)}=B_{s}$, for all $s\in\mathbb{G}(S)$, and consequently $\rho_{\eta}(s)\rho_{\eta}(t)=\rho_{\eta}(st)$, for all $(s,t)\in\mathbb{G}(S)_{2}$.
Applying the inverse functor $\mathbb{S}$ of $\mathbb{G}$ via the ESN Theorem, we obtain an inverse semigroup homomorphism $\mathbb{S}(\rho_{\eta}):S\to\mathcal{I}(A)$. Defining $\beta=(B_{s},\beta_{s})_{s\in S}$ by $\beta_{s}:B_{s^{-1}}\to B_{s}$ with $\beta_{s}(b)=\mathbb{S}(\rho_{\eta})(s)(b)=\eta_{s}(b)$, for all $b\in B_{s^{-1}}$, $s\in S$, we have that $\eta$ being minimal globalization of $\omega$ implies that $\beta$ is the desired globalization of $\alpha$. The uniqueness (up to equivalence) of $\beta$ follows from Theorem 4.7.
Consider the definition of partial skew inverse semigroup ring given in [9], that is, $A\ltimes_{\alpha}S=\mathcal{L}/\mathcal{N}$, where $\alpha$ is a unital partial action of $S$ on a ring $A$, $\mathcal{L}=\bigoplus_{s\in S}A_{s}\delta_{s}$ with component-wise addition and product given by $(a_{s}\delta_{s})(b_{t}\delta_{t})=\alpha_{s}(\alpha_{s^{-1}}(a_{s})b_{t})\delta_{st}$, and $\mathcal{N}=\langle a_{s}\delta_{s}-a_{s}\delta_{t}:s\leq t\text{ and }a_{s}\in A_{s}\rangle$ is an ideal of $\mathcal{L}$. Then the same argument used in Theorem 5.2 also proves the following result.
@@
Let $\alpha$ be a unital partial inverse semigroup action of $S$ on a ring $A$. Assume that $E(S)$ is finite. Let $\beta$ be the minimal globalization of $\alpha$, which is a global action of $S$ on a ring $B$. Then $A\ltimes_{\alpha}S$ and $B\ltimes_{\beta}S$ are Morita equivalent.
Declarations
Author’s Contribution
All authors wrote the main manuscript text and reviewed the manuscript.
Competing Interests
The authors have no competing interests as defined by Springer, or other interests that might be perceived to influence the results and/or discussion reported in this paper.
Ethical Approval
Not applicable.
Availability of Data and Materials
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Funding
There was no funding on the realization of this work.